# CMB delensing with deep learning

### v2

@@ 4.2 Angular power spectrum analysis
$\begin{split}&\Gamma^{XX,lensed}=\text{Mean}(|\frac{C_{\ell}^{XX,lensed}}{C_{\ell}^{XX,unlensed}}-1|),\\ &\Gamma^{XX,delensed}=\text{Mean}(|\frac{C_{\ell}^{XX,delensed}}{C_{\ell}^{XX,unlensed}}-1|),\end{split}$ (19)
where the subscript $XX=\{TT,EE,BB,TE\}$. The subscript $lensed$ corresponds to the quantized CMB lensed power spectrum, while the subscript $delensed$ represents the quantized CMB delensed power spectrum. The term Mean specifically refers to the calculation of the average value of an array that is enclosed within parentheses. The physical interpretation of Equation (19) is that a value of $\Gamma^{XX}$ approaching zero indicates a diminished impact of the lensing effect.
We initiate our investigation with an in-depth analysis of Figure 10, which illustrates the lensed, unlensed, and delensed CMB TT angular power spectrum, as well as their corresponding ratios. Upon careful examination of the image, we observed that the delensed and unlensed angular power spectra are highly similar on larger scales (smaller $\ell$ values). Based on the quantitative index for calculation [Equation (19)], we obtained the results of $\Gamma^{TT,lensed}=0.0194$ and $\Gamma^{TT,delensed}=0.0023$, which indicate the significant impact of the delensing treatment.
Figure 10: The lensed, unlensed, and delensed CMB TT power spectrum, along with their respective ratios. Top panel: The red solid line represents the lensed TT spectrum, the blue solid line represents the unlensed TT spectrum, and the orange solid line represents the delensed TT spectrum; Bottom panel: The pink solid line represents the ratio between the delensed TT spectrum and lensed lensed lensed spectrum, the light blue solid line represents the ratio between the unlensed TT spectrum and the lensed TT spectrum, and the gray solid line represents ${\rm Ratio}=1$.
However, as we focus on smaller scales (corresponding to an increase in $\ell$), the difference between the delensed and unlensed power spectra gradually increases, with the delensed spectrum showing a decreasing trend. This phenomenon reveals the varying performance of the U-Net++ network across different scales. Specifically, as the scale decreases, the delensing effect of U-Net++ diminishes gradually, and the rate of this decrease accelerates. We propose two possible explanations for this finding.
Firstly, we believe that the lensing effect may gradually increase on smaller scales, causing a significant distortion of the original CMB signal. This distortion may exceed the correction capability of the U-Net++ network, leading to a decrease in delensing effectiveness on even smaller scales. Secondly, we considered the impact of the segmentation scheme selection on the results. In this study, we used $N_{side}=4$ as the base resolution for segmentation, which may have adverse effects on the delensing effectiveness on smaller scales.
To investigate the validity of the second hypothesis, we conducted further calculations and analyses. When $N_{side}=4$, each pixel covers approximately $214.86$ square degrees on the sky, corresponding to a multipole moment $\ell$ of around $100$. However, Figure 1 shows significant anomalies only when $\ell$ is greater than $1000$. Hence, we can initially dismiss the segmentation scheme as the primary reason for the decreased delensing effectiveness on smaller scales. Consequently, the most probable reason for the decreased delensed effect of U-Net++ on smaller scales is the first explanation: U-Net++ cannot effectively remove the lensing effect when the lensing effect is substantial.
Figure 11 illustrates the lensed, unlensed and delensed CMB EE angular power spectra, along with their corresponding ratios. Through detailed analysis of the image, we find that the differences among these three power spectra are not significant. Nevertheless, in the ratio plot, a conspicuous oscillation emerges in the power spectrum of the large scale(small $\ell$ values), which are primarily caused by cosmic variance, but have no impact on our research results.
Figure 11: The lensed, unlensed, and delensed CMB EE power spectrum, along with their respective ratios. Top panel: The red solid line represents the lensed EE spectrum, the blue solid line represents the unlensed EE spectrum, and the orange solid line represents the delensed EE spectrum. Bottom panel: The pink solid line represents the ratio between the delensed EE spectrum and lensed spectrum, the light blue solid line represents the ratio between the delensed EE spectrum and the unlensed EE spectrum, and the gray solid line represents ${\rm Ratio}=1$.
To more specifically evaluate the effectiveness of the delensing effect, we also calculated quantitative indicators using Equation (19), obtaining $\Gamma^{EE,lensed}=0.0687$ and $\Gamma^{EE,delensed}=0.0276$. These two values indicate that the EE spectrum has seen a notable improvement after undergoing the delensing treatment, and that the delensing capabilities of the TT and EE spectra are roughly equivalent.
Similar to the case of the TT spectrum, as we focus on the EE spectrum at smaller scales, the difference between the delensed spectrum and the lensed spectrum gradually increases. This is also due to the inherent limitations of the U-Net++ model in terms of its delensing capabilities.
Figure 12 presents the lensed, unlensed, and delensed BB angular power spectra of the CMB, along with their corresponding ratios. Notably, compared to the TT and EE spectra, the lensing potential has a particularly significant impact on the BB spectrum. There are three main reasons for this phenomenon. Firstly, the lensing potential not only converts E-mode signals into B-mode signals, but also reverses the process, converting B-mode signals back into E-mode signals. This may lead to hallucination effects in the U-Net++ network. Secondly, the unlensed B-mode signal itself is extremely weak, even below the level of the conversion signal from E-mode to B-mode caused by the lensing effect. Therefore, under the influence of the lensing effect, the difference between the lensed BB spectrum and the unlensed BB spectrum becomes particularly prominent. This difference is visually reflected in Figure 2, demonstrating the order of magnitude difference in the impact of the lensing effect on the BB spectrum. Finally, the difference between lensed BB and unlensed BB is extremely significant at small scales, approaching $10^{4}$. The processing of such High-Dynamic-Range data may be limited in terms of the effectiveness and generalization capabilities of deep learning.
Figure 12: The lensed, unlensed, and delensed CMB BB power spectrum, along with their respective ratios. Top panel: The red solid line represents the lensed BB spectrum, the blue solid line represents the unlensed BB spectrum, and the orange solid line represents the delensed BB spectrum. Bottom panel: The pink solid line represents the ratio between the delensed BB spectrum and lensed spectrum, the light blue solid line represents the ratio between the delensed BB spectrum and the unlensed BB spectrum, and the gray solid line represents ${\rm Ratio}=1$.
Upon further examination of Figure 12, it can be observed that U-Net++ performs well in delensing on large scales. However, as the scale decreases, its effectiveness gradually diminishes. While this is partly due to the inherent limitations of U-Net++ in delensing, the primary reason lies in the strong influence of the lensing potential on the BB spectrum. This also highlights the deficiency of the generalization ability of deep learning under the U-Net network structure for High-Dynamic-Range data (Makinen et al., 2021; Ni et al., 2022). To quantify this influence, we adopted a similar approach as used for the TT and EE spectra and calculated relevant metrics using Equation (19), namely $\Gamma^{BB,lensed}=2475.83,\Gamma^{BB,delensed}=0.5138$. The results indicate that U-Net++ can effectively perform delensing on large scales; however, on small scales, especially when the multipole moment (ell) is greater than 500, the effectiveness of delensing is almost completely lost. To more intuitively demonstrate these quantitative results, we also performed segmented calculations using Equation (19) for the ranges $5002<\ell<$ 500 and $1000500<\ell<$ 1000. The respective results of $\Gamma^{BB,lensed}_{2<\ell<500}=332.46,\Gamma^{BB,delensed}_{500<\ell<1000}=0.3163$ and $\Gamma^{BB,lensed}_{2<\ell<500}=4614.94,\Gamma^{BB,delensed}_{500<\ell<1000}=0.7110$ further confirm the significant decrease in the effectiveness of U-Net++’s delensing capabilities on small scales.
Figure 13 presents the CMB TE angular power spectra of lensed, unlensed, and delensed cases. Given the characteristics of the TE spectrum, we used the difference in lensing effects as the comparison benchmark. Unlike the auto-correlation power spectra such as TT, EE, and BB, the TE spectrum is a cross-correlation power spectrum that reflects the correlation between the CMB T-mode and E-mode at different scales. From the upper panel of Figure 3, we can clearly observe the anti-peak and peak of the lensed and unlensed $C_{\ell}^{TE}$.
Figure 13: The lensed, unlensed, and delensed CMB TE power spectrum, along with their respective ratios. Top panel: The orange solid line represents the delensed TE spectrum; the red solid line represents the lensed TE spectrum; the blue solid line represents the unlensed TE spectrum. Bottom panel: The light blue solid line demonstrates the differences between the delensed TE spectrum and the unlensed TE spectrum; the pink solid line portrays the differences between the lensed TE spectrum and the unlensed TE spectrum. As a reference benchmark, the grey solid line labeled $\rm Diff=0$ represents a state of perfect consistency where there is no difference between the two spectra.
In the top panel of Figure 13, both the training data (lensed, unlensed) and the prediction data (delensed) exhibit non-zero correlation. Notably, the delensed data also exhibits distinct anti-peak and peak, confirming that our network did not resort to brute-force methods to solve the delensing process but truly learned the underlying differences between lensed and unlensed sky map. However, in the bottom panel of Figure 13, we also notice that the noise level of the prediction data is significantly higher compared to the real data (unlensed), especially at smaller scales. This may be due to the limited utilization of global contextual information by U-Net++, which primarily relies on local information for pixel-level predictions.
5 Conclusions
The continuous advancements in deep learning technology have significantly bolstered its application in cosmological research. These studies not only validate the potential of neural networks as an alternative to traditional methods for correcting CMB lensing effects, but also drive further innovation within the domain. We draw upon this series of research findings and adopt the advanced U-Net++ network architecture to perform refined delensing processing on CMB images, which can accurately identify and extract feature information at the pixel level. This study integrates the sky map segmentation technique, as suggested by researchers like Makinen et al. (2021), into CMB delensing methods for the first time, aimed at overcoming the challenges posed by weak gravitational lensing effects.
In large-scale observations of the CMB, subtle cosmological signals, particularly B-mode polarization signals, are extremely weak and easily contaminated by foreground emissions. Moreover, the inevitable weak gravitational lensing phenomenon distorts the original CMB image during the cosmic evolution, causing mixing conversions between T modes and E modes, as well as between E modes and B modes. Therefore, finding effective means to remove the interference from weak gravitational lensing on CMB data is one of the key challenges in revealing the genuine B-mode polarization signal.
In this work, we employ a novel full sky map segmentation scheme (Makinen et al., 2021; Ni et al., 2022; Gao et al., 2023). While DeepSphere and NNhealpix have made significant advancements in processing full sky MB images, they often require excessively long initialization times, surpassing even model training times when dealing with high-resolution datasets ($N_{side}=2048$). In contrast, our method significantly optimizes this process, ensuring initialization time remains under second-level standard while maintaining efficient training speed. Additionally, to enhance the realism of our simulations resembling actual astronomical observations, we include instrumental noise and effects in our analysis. The primary focus of this research is to investigate whether deep learning methods, particularly the U-Net++ algorithm, can play a pivotal role in tackling the challenge of removing weak gravitational lensing effects from CMB data.
In the delensing study of full sky map for CMB temperature T, polarization Q and U components, the U-Net++ network demonstrates remarkable performance. Evaluating image quality metrics, we find that both SSIM and PSNR show significant improvements. Specifically, when processing T maps, SSIM and PSNR improve by $31.95\%$ and $42.46\%$, respectively, while for Q and U maps, these indices see even more pronounced improvements, increasing approximately by $42\%$ and up to $72\%$, strongly indicating that the U-Net++ network can effectively remove gravitational lensing effects at the pixel level.
At the power spectrum level, we examined the characteristic changes in the TT, EE, BB, and TE spectra. By calculating the quantification function we set, we found that on the delensed TT spectrum, the $\Gamma^{TT,\text{delensed}}$ value was only $0.0023$, significantly lower than the $\Gamma^{TT,\text{lensed}}$ value of $0.0194$ before delensing, indicating a pronounced delensing effect. Similarly, significant improvements were observed in the EE spectrum, with $\Gamma^{EE,\text{delensed}}=0.0276$ and $\Gamma^{EE,\text{lensed}}=0.0687$. Due to the significant differences between the lensed and unlensed BB spectra, the BB delensing effect was even more pronounced, with $\Gamma^{BB,\text{delensed}}=0.5137$ across all scales, compared to $\Gamma^{BB,\text{lensed}}=2473.70$. Although the BB spectrum had already shown notable improvement from its initial state, potential for further enhancement in the delensing effect at smaller scales remained evident. As for the TE spectrum, although the delensing results were significantly affected by noise, the anti-peak and peak features presented in the predicted images suggested that the U-Net++ network had successfully learned the pattern differences between lensed and unlensed spectra and possessed a certain degree of generalization ability. This ruled out the possibility that the model had simply used a brute-force method to fit the delensing results. This implies that the U-Net++ network can understand and effectively simulate the physical process from lensing to delensing.
While the U-Net++ model achieves remarkable results in CMB delensing tasks, we note a decreasing trend in the recovery effect for all components in the small-scale power spectra. Although initially speculated to be due to limitations of the segmentation scheme at extremely high resolutions and small scales, after in-depth exploration of resolution and feature scales, we exclude this possibility.
Based on current analyses, we infer that the U-Net mainly relies on local information for pixel-level predictions, with limited utilization of global context information. For scenarios requiring more accurate segmentation with global context information, additional designs such as self-attention mechanisms or other forms of global pooling operations might be needed to enhance model performance. Consequently, in future work, we will attempt to improve this method by incorporating attention mechanisms, multi-scale fusion, deeper and wider network structure designs, and specific data augmentation strategies.
Looking ahead, our primary focus will be on the delensing of CMB B-mode polarization sky maps. This is because the impact of weak gravitational lensing on CMB B modes is significantly greater compared to T modes and E modes. CMB B-mode polarization sky map’s delensing is a crucial task for detecting primordial gravitational waves and understanding inflationary cosmology. Thus, gravitational lensing effects may potentially mask critical signals of interest, rendering the detection of primordial gravitational waves from cosmic inflation unattainable.
@@ Acknowledgments
We thank Sebastian Belkner for his valuable suggestions and guidance on the correct use of lenspyx. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 11975072, 11835009, and 11875102), the Liaoning Revitalization Talents Program (Grant No. XLYC1905011), the National 111 Project of China (Grant No. B16009), and the Science Research Grants from the China Manned Space Project (Grant No. CMS-CSST-2021-B01).
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.

### v3

@@ IV.2 Angular power spectrum analysis
$\begin{split}\Gamma^{XX}_{\Lambda}={\rm Mean}(|\frac{C_{\ell}^{XX_{unlensed}}}{C_{\ell}^{XX_{delensed}}}-1|)\end{split}$ (19)
where the lower subscript $\Lambda$ represents the QE or the UNet++ algorithm, and the upper subscript $XX=\{{\rm TT},{\rm EE},{\rm BB}\}$. The closer the value of $\Gamma$ is to zero, the better the delensing effect; Conversely, the greater the deviation of $\Gamma$ from zero, the less satisfactory the delensing effect.
Based on the quantitative evaluation metric specified in Equation (19), for the TT spectrum, the error value of the QE algorithm is $\Gamma^{TT}_{Q}E=0.6451$, while that of the UNet++ algorithm is $\Gamma^{TT}_{U}Net++=0.0582$. This indicates that the error of the QE algorithm is approximately 10 times that of the UNet++ algorithm. For the EE spectrum, the error of the QE algorithm is $\Gamma^{EE}_{Q}E=0.4560$, and the error of the UNet++ algorithm is $\Gamma^{EE}_{U}Net++=0.0362$. Here, the error of the QE algorithm is roughly 11 times that of the UNet++ algorithm. In the case of the BB spectrum, the disparity in errors between the two algorithms is even more pronounced. The error of the QE algorithm is $\Gamma^{BB}_{Q}E=0.9224$, and the error of the UNet++ algorithm is $\Gamma^{BB}_{U}Net++=0.0693$. At this point, the error of the QE algorithm is approximately 12 times that of the UNet++ algorithm. Based on these comparisons, it is evident that the performance of UNet++ exhibits a significant improvement over the QE algorithm.
IV.3 Discussion
The results and analysis presented in the previous section demonstrate that the UNet framework achieves remarkably high performance in delensing. This outcome arises from the deliberate design of our proof-of-concept experiment: the network was trained to map noisy, lensed inputs onto idealized B-mode maps that are both noiseless and unlensed. Consequently, the model is encouraged not only to remove distortions induced by gravitational lensing but also to suppress noise, leading to results that appear exceptionally strong.
A further concern is whether such performance merely reflects memorization of the training labels rather than the learning of physically meaningful patterns. To address this, we enforced strict separation between training, validation, and test datasets, and ensured that the cosmological parameters used to generate the test maps differed from those employed in training. The results indicate that the network performs well when the test parameters lie close to the training distribution, whereas its performance degrades for out-of-distribution cases. This distribution-dependent behavior suggests that the model acts analogously to an interpolation scheme: it achieves strong results within the range of the training data but exhibits limited extrapolation capability. Crucially, this behavior confirms that the network has not simply memorized the training labels, but has instead learned nontrivial mappings associated with the gravitational lensing process, effectively ruling out overfitting.
It should be emphasized that our current experiments are entirely based on simulated data. In the process of simulation, certain complexities inevitably present in real observations are omitted, such as non-Gaussian noise, residual foreground contamination, instrumental systematics, and uncertainties related to cosmological models. Therefore, we do not rule out the possibility that the model’s performance may degrade when applied to real observational data. Nevertheless, delensing of the CMB remains a frontier research area. In the absence of large-scale, high-quality observational data, exploring and validating new methods using simulated data is an essential step toward advancing this field.
Looking forward, with the upcoming high-quality observational data from next-generation experiments such as LiteBIRD, CMB-S4, AliCPT, and PIPER, it will become possible to validate and refine these deep learning approaches under conditions closer to real observations. Furthermore, by incorporating more accurate noise modeling, complex foreground components, and instrumental systematics, future studies are expected to systematically evaluate and potentially extend the applicability and robustness of the methods proposed in this work.
V Conclusions
In CMB observations, weak gravitational lensing distorts CMB images, causing mixing between the T-mode and E-mode, as well as between the E-mode and B-mode. Therefore, finding effective methods to remove the effects of weak gravitational lensing on CMB data is one of the key challenges in revealing the true B-mode polarization signal.
We have adopted a relatively novel solution that allows spherical data to be processed using deep learning methods.
Our approach presents several advantages. It can be applied regardless of the resolution of the full-sky map and requires no complex initialization procedures. The data can be used for model training after simple preprocessing of the sky maps. Segmentation errors are inevitably present in sky maps irrespective of the data type but our model retains the original data resolution as much as possible through multiple rotation operations.
During the delensing study of full-sky images of CMB T-mode, polarization E-mode, and B-mode components, the UNet++ network demonstrated outstanding performance. We analyzed it at both the image and angular power spectrum levels.
At the image level, we employed two image quality assessment metrics, SSIM and PSNR. For map patches of T/E/B modes, the UNet++ algorithm showed values of SSIM very close to 1 compared with the QE algorithm, and PSNR values were also within a highly similar range (greater than $30$ dB).
At the power spectrum level, we analyzed the characteristic changes in TT, EE, and BB spectra. By calculating corresponding quantitative indicators, the value of $\Gamma^{TT}_{U}Net++$ was only 0.0582, significantly lower than the delensed $\Gamma^{TT}_{Q}E=0.6451$, which fully demonstrates the significant effectiveness of the delensing method. Similarly, there was a marked improvement observed in the EE spectrum: $\Gamma^{EE}_{U}Net++=0.0362$, while the corresponding $\Gamma^{EE}_{Q}E=0.4560$. The difference between lensed and unlensed BB spectra was particularly pronounced, hence its delensing effect was even more prominent. Specifically, $\Gamma^{BB}_{U}Net++=0.0693$, whereas $\Gamma^{BB}_{Q}E$ approached $1$, further confirming the superior performance of the UNet++ algorithm in this mode. Moreover, by introducing a rotation mechanism, the angular power spectrum improved significantly compared to the non-rotated scenario. However, there is still room for optimizing the delensing effect on small-scale structures.
Furthermore, weak cosmological signals, especially B-mode polarization signals, are particularly susceptible to contamination by foreground radiation. Therefore, in future work, we will consider the impact of foreground components to achieve effective subtraction of CMB foregrounds.
Data Availability
The code used to perform the analysis in this work is publicly available at https://github.com/understars0516/cmb_delensing. The repository includes all necessary scripts, configuration files, and instructions to reproduce the key results and figures presented in this paper.
@@ Acknowledgments
We thank Sebastian Belkner for his valuable suggestions and guidance on the correct use of Lenspyx. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12533001, 12575049, 12473001, 11975072, 11835009, and 11875102), the Liaoning Revitalization Talents Program (Grant No. XLYC1905011), the National 111 Project of China (Grant No. B16009), and the Science Research Grants from the China Manned Space Project (Grant No. CMS-CSST-2025-A02).